OpenStudy (leozap1):

If I had a positive base with a negative exponent that would make it positive correct?

OpenStudy (gorica):

yes

OpenStudy (leozap1):

Thanks

OpenStudy (leozap1):

No it would be negative not positive -_-

OpenStudy (anonymous):

it really all depends on the number the exponent is like if it were -1 you d multiply the number nby itself one time and it would be negetive

OpenStudy (anonymous):

but like i said it all depends on the exponents number.

OpenStudy (leozap1):

Ok ty

OpenStudy (anonymous):

ur welcome im happy to help

OpenStudy (ivettef365):

the negative exponents just makes the equation a fraction

OpenStudy (ivettef365):

if you have a^-1 it is 1/a

OpenStudy (ivettef365):

now if the base is negative and the exponent is positive, then the result depends on what number is on the exponent.

OpenStudy (leozap1):

I have the problem -5 * 1/9 after I work out the exponit would it equal -(1/45)?

OpenStudy (ivettef365):

what is the exponent on that equation ?

OpenStudy (leozap1):

I am doing a exponential function the problem would be -5 * (3^-2)

OpenStudy (ivettef365):

ok 3^-2 = 1/9 -5 . 1/9 = -5/9

OpenStudy (leozap1):

When I place that on a graph how would I though?

OpenStudy (ivettef365):

so you are not only solving, you are graphing. what is the exact question

OpenStudy (anonymous):

Well usually negative *s positive is negative ..

OpenStudy (leozap1):

I am working them all out to fill in the chart